OpenStudy (anonymous):

find the sum 4/16 + 2/13

OpenStudy (ybarrap):

look for a common denominator. Easiest way, just multiply 16*13, that's common.

OpenStudy (ybarrap):

Then you do this: 4*13/(16*13) + 2*16)/16*13) = (4*13+2*15)/(16*13)

OpenStudy (ybarrap):

See the pattern?

OpenStudy (anonymous):

no

OpenStudy (ybarrap):

why did I multiply 4*13 in the first term?

OpenStudy (ybarrap):

Look at the denominator for that term, there is also a 13 that cancels out the 13 in the numerator.

OpenStudy (ybarrap):

We did this step so that we can combine numerators easily, like this (4*13+2*15)

OpenStudy (ybarrap):

Notice that 4*13/(16*13) is still just 4/16 like in your original equation (1st term)

OpenStudy (ybarrap):

I just put a 13/13 term there, this is just 1.

OpenStudy (anonymous):

ok? im sorry i dont get this so it takes me a long time to get it

OpenStudy (ybarrap):

bottom line, you need a common denominator to add numerators like I did. Find the common denominator, and you have it made

OpenStudy (anonymous):

oh well

OpenStudy (ybarrap):

it like if you had 1/2 a pie and 1/3 of a pie, what do you get when you add them? You need a way to compare 1/2 and 1/3. A common denominator lets you do that

OpenStudy (ybarrap):

1/2 a pie is the same as 3/6 of a pie and 1/3 of a pie is the same as 2/6 of a pie. Now you know that 3/6 + 2/6 = (3 + 2)/6 = 5/6 of a pie, almost a whole pie!

OpenStudy (anonymous):

ok